The motor of a overhead crane raises a load of 810kg at a velocity of 0.4m/s. the efficiency is 80 percent.

What is the output power?


1
Expert's answer
2021-04-28T07:42:32-0400

Force experienced by crane, "F=mg=810\times9.8=7938\space N"

Velocity of load, "v=0.4\space m/s"

Input Power "=F\times v=3175.2\space W"

Efficiency of crane, "\eta=80\%"

Output Power "=\dfrac{Input\space Power}{\eta}=\dfrac{3175.2}{0.80}=3969 \space W"